# Automatically Advancing LLM Expertise in Technology Judgment

### v3

@@ Appendix F Results for Majority Voting Schema
F.1 Performances of other LLMs with varying sizes on splitting patents
Figure 7: Ask LLMs "Do you think they are the same patent?", given that they are actually different. Judge using majority voting, and 1 = same. Models are instruction-tuning versions.
Figure 8: Ask LLMs "Do you think they are the same patent?", given that they are actually different. Judge using majority voting, and 1 = same. Models except for GPT-4o-mini are instruction-tuning versions.
Figure 9: Ask LLMs "Do you think they are different patents?", given that they are actually different. Judge using majority voting, and 1 = yes. Models except for GPT-4o-mini are instruction-tuning versions.
F.2 LLMs do not know what they have known
Figure 10: Asking Llama 3.1 8B instruct "Do you think they are the same patent?" on actually different patents (1 = wrong answer "yes"). LLM answers using majority voting. It does not know what it has known. These results are statistically significant: Self-talk still helps improve the baseline judge but not as good as scientific self-questioning: $q_{1}$ to $q_{5}$ p=0.000; $q_{6}$ p=0.006.
F.3 Smaller models’ questions help mid-size models more than those from larger models
@@
Appendix G Discussion of Limitations and Future Work
Several limitations and directions for future work remain.

### v4

@@ Appendix F Results for Majority Voting Schema
Figure 6: Ask LLMs "Do you think they are the same patent?", given that they are actually different. The judge is using majority voting, and 1 = same (wrong answer). Llama models are instruction-tuning versions.
Figure 7: Ask LLMs "Do you think they are different patents?", given that they are actually different. The judge is using majority voting, and 1 = yes (correct). Llama models are instruction-tuning versions.
Figure 8: Asking Llama 3.1 8B instruct "Do you think they are the same patent?" on actually different patents (1 = wrong answer "yes"). LLM judges patents using majority voting. These results are statistically significant: Self-talk still helps improve the baseline judge but not as good as scientific answers: $q_{1}$ to $q_{5}$ p=0.000; $q_{6}$ p=0.006.
@@
Appendix G Pretraining Exposure
The key to evaluating the model understanding is that the tests we created are unlikely to be in the training data. If they are, it might be that the models simply regrgitate case that they have already seen before and not demonstrate understanding. Note that we are less concerned with whether the patent texts alone are present in the pretraining data, because we deliberately construct a pairing mechanism whose likelihood of having been encountered during pretraining is extremely low. The patent set used in the paper contains roughly $n=1.5\times 10^{6}$ patent texts. Each text averages about $400$ words, or approximately $520$ byte‑pair‑encoded (BPE) tokens [^8] Rule of thumb: one English word $\approx 1.3$ BPE tokens. See the OpenAI Help Center article “What are tokens and how to count them?” https://help.openai.com/en/articles/4936856-what-are-tokens-and-how-to-count-them.. The number of unordered pairs is $\tfrac{n(n-1)}{2}$, so concatenating every possible pair would require approximately $1.2\times 10^{15}\ \text{tokens}$. That total is multiple orders of magnitude larger than some of the largest publicly reported training corpora, even though a large share of these corpora is (unrelated) github code, books, Wikipedia, etc: GPT models were trained on $\sim 3\times 10^{11}$ tokens [^9] Brown, T.B. et al.(2020), “Language Models are Few‑Shot Learners.” and Llama‑70B on $\sim 2\times 10^{12}$ tokens [^10] Touvron, H. et al.(2023), “Llama: Open Foundation and Fine‑Tuned LLMs.”. The size of Redpajama English tokens is $\sim 2\times 10^{13}$  [^11] https://github.com/togethercomputer/RedPajama-Data.
Appendix H Example Questions
Qualitative inspection of the generated questions suggests that answering them should not require substantial novel information beyond the models’ training (echoing the missing vs. unused knowledge hypothesis). Many questions are straightforward and focus on established concepts or systems. For example, questions involving the TCP Port Service Multiplexer (TCPMUX) relate to long-standing infrastructures, with patents contributing only incremental advancements. As shown in our original paper, 68% of the Llama 1B model’s questions and 44% of the Llama 405B model’s questions include phrases like “what is” or “what are,” indicating the dominance of factual, definitional queries. A sample of questions can be seen in Figure 10.
Figure 10: Example questions.
Appendix I Ablation Studies
Figure 11: More scientific-QA improves the judge over the placebo answer: For panel (a): $p_{q1}$ =0.0868, $p_{q2}$ =0.9270, $p_{q3}$ =0.0934, $p_{q4}$ =0.0000, $p_{q5}$ =0.0000, $p_{q6}$ =0.0021. We ask "are they the same?" ("yes" is the wrong answer) given the human-rated ground truth of being different patents.
We gradually add scientific QA pairs from 1-6 for each patent to the original pair of patent texts, and then have Llama 3.1 8B re-judge. As shown in Figure 11, the baseline represents judge accuracy and confidence without any self-questioning or scientific retrieval. The placebo condition uses the same question but adds a randomly selected piece of text from the same paper as the found answers, matched in length to the actual answers. This is aiming for ablation studies on retrieval quality – what if the retrieved answer is of very low quality (even unrelated)? We observe that as more QAs from science are added, judge accuracy significantly increases and becomes better than placebo answers. The confidence of both correct and incorrect answers decreases, but the drop is significantly greater for incorrect answers. Interestingly, even the placebo answer — though merely a random, same-length excerpt from the paper — enhances understanding to some extent. However, unlike actual scientific answers, its effect does not scale with the number of questions, meaning there is some initial signal that helps judgment but LLM does not learn more over more rounds of relatively unrelated science. This echoes with the findings in the main paper: (1) the question alone already facilitates understanding, functioning as cues for better deployment of internal knowledge (we explored this further in Section 5.1), and (2) the placebo text still contains scientific content, which can contribute to comprehension.